Consider the following reaction in an open flask:


A 155.0 g sample of CaCO3 is placed in the flask and HCl is added. The reaction consumes HCl at an average rate of 0.200mol/min for 10.0 min. What mass of CaCO3 remains?


CaCO3 + 2HCl = CaCl2 + H2O + CO2




A 155.0 sample of CaCO3 is placed in the flask and HCl is added. The reaction consumes HCl at an average rate of 0.200 mol/min for 10.0 min. What mass of CaCO3 remains? (molar mass Ca=40, C=12, O=16).


1
Expert's answer
2021-06-23T05:17:30-0400

n (HCl)reacted = 0.200 mol/min x 10 min = 2.00 mol

n (CaCO3)reacted =2.00 mol HCl x (1 mol CaCO3 / 2 mol HCl) = 1 mol

g (CaCO3)reacted = 1 mol x 100.1 g/mol = 100.1 g

n (CaCO3)left = 155.0 - 100.1 = 54.9 g